OpenStudy (anonymous):

the length of a tabletop is 5m greater than the width. the area is 84m^2 find dimensions of the table

OpenStudy (anonymous):

x(x+5) =84 x^2 + 5x -84 =0 (x-7)(x+12) x=7, x=-12 negative 12 is eliminated so the width =7 and length = 12 is that right

OpenStudy (ajprincess):

u r right.:)